OpenStudy (anonymous):

2Pb(s)+O2(g)--->2PbO(s) In the following reaction, 451.4 g of lead reacts with excess oxygen forming 349.7 g of lead(II) oxide. Calculate the percent yield of the reaction.

OpenStudy (aaronq):

find the theoretical yield then use: \(percent ~yield=\dfrac{theoretical-actual}{theoretical}*100\%\)

OpenStudy (anonymous):

which is which? :/ I'm sorry I'm not good at this

OpenStudy (aaronq):

actual is what is given in the question. theoretical is the most you could make - found through stoichiometry

OpenStudy (anonymous):

oh okay how do you find the most you could make?

OpenStudy (aaronq):

Convert the mass of reactant to moles \(\sf moles=\dfrac{mass}{Molar~mass}\) Next, use the stoichiometric coefficients and the moles to find moles of product \(\sf \dfrac{moles~of~Pb}{Pb's~coefficient}=\dfrac{moles~of~PbO}{PbO's~coefficient}\) Convert moles to mass (using the first formula) use the perfect yield formula

OpenStudy (korosh23):

@aaronq If actual yield is not mentioned, is there any way to find percent yield?

OpenStudy (korosh23):

For instance imagine if we have only theoretical mass and not actual mass and the question has asked us to find the percent yield?

OpenStudy (korosh23):

@aaronq Let me know.

OpenStudy (korosh23):

@lindsh Do you know the actual yield inside your question?

OpenStudy (anonymous):

The actual yield is supposedly given in the problem but I'm not sure. that's why I'm here lol

OpenStudy (korosh23):

Yes that is confusing because I am doing the same stuff. The actual yield must be give to find percent yield. Other than that it is almost impossible.

OpenStudy (korosh23):

Do not close the question. Let him come back and answer my question. Thx

OpenStudy (anonymous):

alright @aaronq I think I got the wrong answer..

OpenStudy (aaronq):

if the actual yield is not given there is not way that you could find that out, so no you can't do the question.

OpenStudy (aaronq):

@korosh23